suppose an item is manufactured by three mavhines X,Y and Z.all the three machines have equal capacity and operated at the same rate.it is known that the percentage of the defective items produced by X,Y and Z are 2,7 and 12 percent respectively.all the items produced by X,Y and Z are put into one bin.from yhis bin,one item is drawnat random is found to bedefective.what is the probabilty that this item is produced by y?

Grade:

2 Answers

Chetan Mandayam Nayakar
12 years ago

answer=((1/3)*7)/(((1/3)*2)+((1/3)*7)+((1/3)*12))

ashish jangid
12 years ago

answer  

p(item produced by y)= no of defective items by y / total defective items

   p(E)= 7/2+7+12= 7/21=0.333